OpenStudy (deadshot):

What is the exact value of \[\tan \frac{ \pi }{ 8 }\] ?

OpenStudy (luigi0210):

Wouldn't you use the half angle formula or something along those lines?

OpenStudy (deadshot):

I honestly have no idea of what I should do to start solving it.

OpenStudy (luigi0210):

Look up half-angle formulas/

OpenStudy (deadshot):

I found this, but I'm still confused. http://mathworld.wolfram.com/Half-AngleFormulas.html

OpenStudy (luigi0210):

Yes, there you go. Now what you have to do is use the one that involves tan

OpenStudy (deadshot):

Okay

OpenStudy (luigi0210):

And if you look at this, you can see that there are 3 ways to find the half-angle of tan: http://www.wyzant.com/resources/lessons/math/trigonometry/half-angle-double-angle-formulas

OpenStudy (anonymous):

tan x/2 = cosecx - cotx Let pi/4 = x, cosec pi/4 - cot pi/4 = rt2 - 1

OpenStudy (luigi0210):

So let's use: \[\LARGE tan\frac{x}{2}=\frac{1-cosx}{sinx}\] Now, what radian do you know, when divided by 2 gives you \[\LARGE \frac{pi}{8}\]

OpenStudy (anonymous):

Let tan(pi/8) = y and tan2x = 2 tanx / (1 - tan^2 x) so tan (pi/4) = 2 tan (pi/8) / ( 1 - tan^2 (pi/8)) 1 = 2y/(1-y^2) 1-y^2 = 2y y^2 + 2y - 1 = 0 y = -1 +/- sqrt(2) since tan (pi/8) is in first quadrant, the value should be positive. Therefore the answer is -1 + sqrt(2)

OpenStudy (deadshot):

Thanks!

OpenStudy (anonymous):

exact value till 5 decimal space 0.41421